# Are “Solved Issues” in SWE-bench Really Solved Correctly? An Empirical Study

### v1

@@ Abstract.
Automated issue solving aims to resolve real-world issues in software repositories. The most popular benchmarks for automated issue solving are SWE-bench and its human-filtered subset SWE-bench Verified, which are widely used to evaluate foundation models and software engineering agents. These benchmarks leverage testing to validate generated patches. However, because testing is rarely exhaustive, a patch may pass the tests but nevertheless fail to match the developers’ expectations. Unfortunately, it is currently unclear to what extent evaluations performed with SWE-bench suffer from such plausible but incorrect patches. This paper presents an in-depth empirical study of the correctness of plausible patches generated by three state-of-the-art issue-solving tools (CodeStory, LearnByInteract, and OpenHands) evaluated on SWE-bench Verified. We extensively test and inspect generated patches, and compare them against human-written ground truth patches. The core of our methodology is a novel technique PatchDiff for differential patch testing, which automatically exposes behavioral discrepancies between two patches. Our findings reveal critical weaknesses in SWE-bench’s patch validation mechanism, which causes 7.8% of all patches to count as “correct” while failing the developer-written test suite. Moreover, our novel automated technique reveals that even more (29.6%) plausible patches induce different behavior than the ground truth patches. These behavioral differences are often due to similar, but divergent implementations (46.8%) and due to generated patches that adapt more behavior than the ground truth patches (27.3%). Our manual inspection shows that 28.6% of behaviorally divergent patches are certainly incorrect. Combined, the different weaknesses lead to an inflation of reported resolution rates by 6.2 absolute percent points. Our findings are a call to arms for more robust and reliable evaluation of issue-solving tools. We envision our automated differential patch testing technique to be useful for this purpose.
copyright: acmlicensedjournalyear: 2018doi: XXXXXXX.XXXXXXXisbn: 978-1-4503-XXXX-X/2018/06copyright: none
@@ 1. Introduction
Automated issue solving aims to address real-world issues in software repositories, and holds great potential to reduce maintenance costs and improve software quality. SWE-bench (Jimenez et al., 2023) stands out as the most popular benchmark for automated issue solving, comprising 2,294 tasks from 12 well-maintained Python repositories. Each task is provided with an issue statement describing the specific task to accomplish and the repository version where the issue is to be solved. The tool to be evaluated is then required to generate a patch to address the issue. SWE-bench evaluates the generated patch by running tests associated with the issue, including at least one fail-to-pass test to confirm the issue is resolved, and, where available, several pass-to-pass tests to ensure no regression in existing functionality. Recently, OpenAI hired 93 developers to manually identify and filter out those SWE-bench tasks with overly specific and even issue-unrelated tests, and curated a subset comprising 500 tasks, known as SWE-bench Verified (swe, [n. d.]). The SWE-bench leaderboard is getting immense attention. It has been used to judge the merit of new techniques proposed in academic papers (Zhang et al., 2024; Xia et al., 2025), and also to evaluate commercial tools for assessing their potential value (wan, [n. d.]; iso, [n. d.]; bla, [n. d.]). Moreover, SWE-bench Verified has been widely used to evaluate the coding abilities of state-of-the-art foundation models, such as OpenAI GPT-o1 (gpt, [n. d.]) and Anthropic Claude-3.5 (cla, [n. d.]).
However, due to practical limitations, test suites are rarely exhaustive and often suffer from weaknesses (Smith et al., 2015). Following the literature on automated program repair (Liu et al., 2019), we refer to patches that pass their corresponding validation process as plausible patches. Validation with weak test suites can result in plausible but incorrect patches, inflating the performance of the evaluated tools, and possibly leading to incorrect conclusions about their abilities. Although others have noticed the problem of weak test suites in SWE-bench (swe, [n. d.]; Aleithan et al., 2024), it is unclear to what extent evaluations performed with SWE-bench suffer from this problem, and there is a notable lack of effective methodologies for detecting plausible but incorrect patches.
This paper conducts an in-depth empirical study of the correctness of plausible generated patches on SWE-Bench. We focus on the high-quality, human-filtered, and widely used subset SWE-bench Verified, and conduct this study on the plausible patches generated by the state-of-the-art issue-solving tools, i.e., CodeStory (cod, [n. d.]), LearnByInteract (Su et al., 2025), and OpenHands (Wang et al., 2024a). Our study addresses four research questions:
@@
To enable this study, we present a novel differential patch testing technique named PatchDiff, which aims to generate test cases that expose behavioral discrepancies between the plausible patch and the developer-written oracle patch. We refer to such generated tests as differentiating tests. PatchDiff leverages Large Language Models (LLMs) for test generation and is enabled by a call-trace-based method to identify appropriate target functions and construct useful contextual information for LLMs. Specifically, we first leverage PatchDiff to generate differentiating tests for each plausible patch and pick out behaviorally divergent patches, and then manually assess the correctness of these suspicious patches with the help of differentiating tests. This approach provides concrete evidence of invalid functionality, enables more focused and objective patch validation, and reduces manual validation effort.
@@
  • Replication package. A replication package (pac, [n. d.]) including the implementation of PatchDiff and the results of our study.
@@ 2.1. SWE-bench and SWE-bench Verified
SWE-bench (Jimenez et al., 2023) is the most popular benchmark for automated issue solving. As described in Section 1, to assess the effectiveness of an issue-solving tool on SWE-bench, the user needs to generate patches for each task based on the corresponding issue statement and the buggy repository version. SWE-bench also provides a test patch for each task, which includes the changes made to test files in the PR that resolves the corresponding issue. The full set of SWE-bench is shown to consist of low-quality and noise instances (swe, [n. d.]). OpenAI curated a high-quality subset known as SWE-bench Verified, which consists of 500 samples.
To validate the correctness of a generated patch, both SWE-bench and SWE-Bench Verified run the test files modified in the test patch after applying both the test patch and the generated patch to the buggy repository version. If these test files all pass, the patch is regarded as correct (har, [n. d.]). Although such test files are likely to be relevant to the issue, they do not necessarily cover all the functionalities that can be affected by the generated patches. As a result, this validation process is weak and may accept incorrect patches that violate the uncovered functionalities.
@@ 3.1. Issue Solving Tools and Patches
Our empirical study focuses on SWE-bench Verified, a high-quality, human-validated subset of SWE-bench. The study is conducted on the plausible patches generated by three state-of-the-art issue-solving tools, i.e., CodeStory Midwit Agent + swe-search (cod, [n. d.]) (hereinafter, CodeStory), Learn-by-interact (Su et al., 2025) (hereinafter, LearnByInteract), and OpenHands + CodeAct v2.1 (Wang et al., 2024a) (hereinafter, OpenHands), on SWE-Bench Verified. These tools are selected because they open source their implementations and/or release their papers or technical reports, making them suitable for further research. Note that for each tool and each issue in SWE-bench Verified, there is at most one plausible patch.
@@ 3.2. Differential Patch Testing
To enable this study, we proposes an automated differential patch testing technique named PatchDiff. Given an issue-solving task providing the issue statement and the buggy repository version, its test patch $P_{t}$, its oracle patch $P_{o}$, and a patch $P_{g}$ generated for it, PatchDiff leverages an LLM to generate a set of tests that can reveal the behavioral differences between $P_{g}$ and $P_{g}$, i.e., differentiating tests. We refer to the repository version with only $P_{t}$ applied as $R_{t}$ and with both $P_{t}$ and $P_{g}$ or $P_{o}$ applied as $R_{g}$ or $R_{o}$, respectively. PatchDiff first checks whether $P_{g}$ and $P_{o}$ are syntactically identical without considering comments, and omits identical $P_{g}$. Then, PatchDiff leverages a call-trace-based method to identify appropriate target functions and extract contextual code. Finally, PatchDiff prompts an LLM to generate and repair tests for target functions, and filters out unqualified tests.
@@ 3.2.1. Target Function Identification
To generate tests, PatchDiff first needs to determine the target function to be tested. It regards a function that satisfies the following criteria as a target function: (1) The function is a patch-modified function or a function that directly or indirectly invokes a patch-modified function. (2) The function is not defined within test files. (3) The function is directly invoked by developer-written tests. Such a function is related to patches and not related to tests, and developers usually have specific expectations regarding its behavior. To identify target functions, we first instruct each patch-modified function in $R_{g}$ and $R_{o}$, and run all test files in $R_{g}$ and $R_{o}$ to collect call traces. Each call trace starts from a test function in $R_{g}$ or $R_{o}$ and ends at a patch-modified function. The first non-test function in each call trace satisfies the criterion mentioned above and PatchDiff annotates it as a target function. Different call traces can have the same target function. So a target function may correspond to multiple call traces.
@@ 3.2.2. Contextual Code Extraction
In each call trace, the functions in test files provide information about how to invoke the target function, and the functions in non-test files are either the target function or illustrate how the target function utilizes the patch-modified function. So all the functions in the call trace are useful for generating differentiating tests. For each target function, we merge the functions in all of its call traces. For each of these functions, we further map it to its before-patch version in $R_{t}$ based on its class and function names. The mapped functions provide the contextual information before patches are applied and are referred to as context functions. Only providing function definitions to the LLM may miss critical contextual information such as class information. Therefore, we extend context functions to construct contextual code. Specifically, we collect all Python files that contain the target function or at least one context/patch-modified function from $R_{t}$. We remove the functions that are not target, context, or patch-modified functions from these files. Any classes rendered empty after this deletion are also discarded. In addition, in test files, we annotate code lines where the target function is invoked with a comment to direct the LLM’s attention. All the streamlined files form the contextual code of this target function.
@@ 3.2.4. Unqualified Test Filtering
PatchDiff aims at generating tests for target functions to expose meaningful behavioral discrepancies. However, since LLMs do not always follow the instructions, the generated tests may not only examine the specified target function. To increase the probability of exposing meaningful behavioral discrepancies, PatchDiff filters out such tests. Specifically, we instrument patch-modified functions and execute each generated differentiating test to collect call traces. For each differentiating test, if there is a call trace where the first production function is not a target function, we filter it out. Because this test does not only examine target functions. We further filter out flaky tests. In detail, for each remaining differentiating test, we run it under $P_{g}$ and $P_{o}$ for 20 times each. If it passes under one patch for all 20 times and fails under the other patch for at least one time, we regard it as valid. Otherwise, we filter out it. All the remaining differentiating tests are regarded as the output of PatchDiff.
@@ 4.2.2. Comparison with Existing Test Generators
Approach: We use two representative test generators for Python: Pynguin (Lukasczyk and Fraser, 2022) and CoverUp (Pizzorno and Berger, 2024). Pynguin is the state-of-the-art search-based test generation tool for Python. CoverUp is a representative LLM-based test generation tool with its implementation publicly available and easy to use. Please note that both Pynguin and CoverUp support only Python 3.10 or higher, a requirement met by only 14.8% of the 500 tasks in SWE-bench Verified. Consequently, our evaluation is confined to the plausible patches generated by the three tools that are under a compatible Python version, yielding 133 plausible patches. For both test generators, we designate the patch-modified files as the target modules, and tests are generated on the repository with the oracle patch applied. Following PatchDiff’s setting, we also use gpt-4o-mini-2024-07-18 as the underlying LLM of CoverUp for a fair comparison. Since CoverUp supports only PyTest and fails on the instances employing other test frameworks (e.g., Django’s customized framework), we do not provide original tests to CoverUp to ensure it can handle the 133 instances.
@@
Results: Table  3 presents the evaluation results. Pynguin produces at least one test file for only 3 plausible patches because it frequently encounters exceptions during generation. For example, for 93 patches, Pynguin raises an AttributeError which states that in the file pynguin/testcase/execution.py, attribute IN cannot be found in enum Compare, which may be a bug in Pynguin. CoverUp produces at least one test file only for 40 patches due to the difficulties it has in constructing appropriate execution environments (such as the database setup required by Django). None of the tests generated by Pynguin or CoverUp reveal behavioral discrepancies between the plausible and oracle patches, i.e., none of these tests are differentiating tests. In contrast, PatchDiff produces at least one test file for 117 patches and successfully generates differentiating tests for 56 of them. Note that among the 133 plausible patches, PatchDiff only attempts to generate tests for 117 of them. 15 of them are identical to their corresponding oracle patches and are omitted by PatchDiff, and 1 patch contains syntax errors and cannot be correctly parsed by the unidiff library (uni, [n. d.]). These results underscore the effectiveness and necessity of PatchDiff in generating differentiating tests.
Answers to RQ2: PatchDiff generates differentiating tests for 29.6% of the plausible patches. This suggests that a significant proportion of plausible patches behave differently from the oracle patches. In contrast, regression test generation baselines cannot generate differentiating tests for any patch.

### v2

@@ Abstract.
Automated issue solving aims to resolve real-world issues in software repositories. The most popular benchmarks for automated issue solving are SWE-bench and its human-filtered subset SWE-bench Verified, which are widely used to evaluate foundation models and software engineering agents. These benchmarks leverage testing to validate generated patches. However, because testing is rarely exhaustive, a patch may pass the tests but nevertheless fail to match the developers’ expectations. Unfortunately, it is currently unclear to what extent evaluations performed with SWE-bench suffer from such plausible but incorrect patches. This paper presents an in-depth empirical study of the correctness of plausible patches generated by three state-of-the-art issue-solving tools (CodeStory, LearnByInteract, and OpenHands) evaluated on SWE-bench Verified. We extensively test and inspect generated patches, and compare them against human-written ground truth patches. The core of our methodology is a novel technique for differential patch testing, called PatchDiff, which automatically exposes behavioral discrepancies between two patches. Our findings reveal critical weaknesses in SWE-bench’s patch validation mechanism, which causes 7.8% of all patches to count as “correct” while failing the developer-written test suite. Moreover, our novel automated technique reveals that even more (29.6%) plausible patches induce different behavior than the ground truth patches. These behavioral differences are often due to similar, but divergent implementations (46.8%) and due to generated patches that adapt more behavior than the ground truth patches (27.3%). Our manual inspection shows that 28.6% of behaviorally divergent patches are certainly incorrect. Combined, the different weaknesses lead to an inflation of reported resolution rates by 6.4 absolute percent points. Our findings are a call to arms for more robust and reliable evaluation of issue-solving tools. We envision our automated differential patch testing technique to be useful for this purpose.
Automated Issue Solving, Differential Testing, Patch Evaluation
journalyear: 2026copyright: rightsretainedconference: 2026 IEEE/ACM 48th International Conference on Software Engineering; April 12–18, 2026; Rio de Janeiro, Brazilbooktitle: 2026 IEEE/ACM 48th International Conference on Software Engineering (ICSE ’26), April 12–18, 2026, Rio de Janeiro, Brazildoi: 10.1145/3744916.3764576isbn: 979-8-4007-2025-3/26/04ccs: Software and its engineering
@@ 1. Introduction
Automated issue solving aims to address real-world issues in software repositories, and holds great potential to reduce maintenance costs and improve software quality. SWE-bench (Jimenez et al., 2023) stands out as the most popular benchmark for automated issue solving, comprising 2,294 tasks from 12 well-maintained Python repositories. Each task is provided with an issue statement describing the specific task to accomplish and the repository version where the issue is to be solved. The tool to be evaluated is then required to generate a patch to address the issue. SWE-bench evaluates the generated patch by running tests associated with the issue, including at least one fail-to-pass test to confirm the issue is resolved, and, where available, several pass-to-pass tests to ensure no regression in existing functionality. Recently, OpenAI hired 93 developers to manually identify and filter out those SWE-bench tasks with overly specific and even issue-unrelated tests, and curated a subset comprising 500 tasks, known as SWE-bench Verified (swe, 2024). The SWE-bench leaderboard is getting immense attention. It has been used to judge the merit of new techniques proposed in academic papers (Zhang et al., 2024; Xia et al., 2025), and also to evaluate commercial tools for assessing their potential value (wan, 2025; iso, 2025; bla, 2025). Moreover, SWE-bench Verified has been widely used to evaluate the coding abilities of state-of-the-art foundation models, such as OpenAI GPT-o1 (gpt, 2025) and Anthropic Claude-3.5 (cla, 2024).
However, due to practical limitations, test suites are rarely exhaustive and often suffer from weaknesses (Smith et al., 2015). Following the literature on automated program repair (Liu et al., 2019), we refer to patches that pass their corresponding validation process as plausible patches. Validation with weak test suites can result in plausible but incorrect patches, inflating the performance of the evaluated tools, and possibly leading to incorrect conclusions about their abilities. Although others have noticed the problem of weak test suites in SWE-bench (swe, 2024; Aleithan et al., 2024), it is unclear to what extent evaluations performed with SWE-bench suffer from this problem, and there is a notable lack of effective methodologies for detecting plausible but incorrect patches.
This paper conducts an in-depth empirical study of the correctness of plausible generated patches on SWE-bench. We focus on the high-quality, human-filtered, and widely used subset SWE-bench Verified, and conduct this study on the plausible patches generated by the state-of-the-art issue-solving tools, i.e., CodeStory (cod, 2025), LearnByInteract (Su et al., 2025), and OpenHands (Wang et al., 2024a). Our study addresses four research questions:
@@
To enable this study, we present a novel differential patch testing technique, named PatchDiff, which aims to generate test cases that expose behavioral discrepancies between the plausible patch and the developer-written oracle patch. We refer to such generated tests as differentiating tests. PatchDiff leverages Large Language Models (LLMs) for test generation and is enabled by a call-trace-based method to identify appropriate target functions and construct useful contextual information for LLMs. Specifically, we first leverage PatchDiff to generate differentiating tests for each plausible patch and identify behaviorally divergent patches, and then manually assess the correctness of these suspicious patches with the help of the differentiating tests. This approach provides concrete evidence of invalid functionality, enables more focused and objective patch validation, and reduces manual validation effort.
@@
  • Replication package. A replication package (pac, 2025) including the implementation of PatchDiff and the results of our study.
@@ 2.1. SWE-bench and SWE-bench Verified
SWE-bench (Jimenez et al., 2023) is the most popular benchmark for automated issue solving. As described in Section 1, to assess the effectiveness of an issue-solving tool on SWE-bench, the user needs to generate patches for each task based on the corresponding issue statement and the buggy repository version. SWE-bench also provides a test patch for each task, which includes the changes made to test files in the PR that resolves the corresponding issue. The full set of SWE-bench is shown to consist of low-quality and noise instances (swe, 2024). OpenAI curated a high-quality subset known as SWE-bench Verified, which consists of 500 samples.
To validate the correctness of a generated patch, both SWE-bench and SWE-bench Verified run the test files modified in the test patch after applying both the test patch and the generated patch to the buggy repository version. If these test files all pass, the patch is regarded as correct (har, 2025). Although such test files are likely to be relevant to the issue, they do not necessarily cover all the functionalities that can be affected by the generated patches. As a result, this validation process is weak and may accept incorrect patches that violate the uncovered functionalities.
@@ 3.1. Issue Solving Tools and Patches
Our empirical study focuses on SWE-bench Verified, a high-quality, human-validated subset of SWE-bench. The study is conducted on the plausible patches generated by three state-of-the-art issue-solving tools, i.e., CodeStory Midwit Agent + swe-search (cod, 2025) (hereinafter, CodeStory), Learn-by-interact (Su et al., 2025) (hereinafter, LearnByInteract), and OpenHands + CodeAct v2.1 (Wang et al., 2024a) (hereinafter, OpenHands), on SWE-bench Verified. These tools are selected because they open source their implementations and/or release their papers or technical reports, making them suitable for further research. Note that for each tool and each issue in SWE-bench Verified, there is at most one plausible patch.
@@ 3.2. Differential Patch Testing
To enable this study, we proposes an automated differential patch testing technique named PatchDiff. Given an issue-solving task providing the issue statement and the buggy repository version, its test patch $P_{t}$, its oracle patch $P_{o}$, and a patch $P_{g}$ generated for it, PatchDiff leverages an LLM to generate a set of tests that can reveal the behavioral differences between $P_{g}$ and $P_{o}$, i.e., differentiating tests. We refer to the repository version with only $P_{t}$ applied as $R_{t}$ and with both $P_{t}$ and $P_{g}$ or $P_{o}$ applied as $R_{g}$ or $R_{o}$, respectively. PatchDiff first checks whether $P_{g}$ and $P_{o}$ are syntactically identical without considering comments, and omits identical $P_{g}$. Then, PatchDiff leverages a call-trace-based method to identify appropriate target functions and extract contextual code. Finally, PatchDiff prompts an LLM to generate and repair tests for target functions, and filters out unqualified tests.
@@ 3.2.1. Target Function Identification
To generate tests, PatchDiff first needs to determine the target function to be tested. It regards a function that satisfies the following criteria as a target function: (1) The function is a patch-modified function or a function that directly or indirectly invokes a patch-modified function. (2) The function is not defined within test files. (3) The function is directly invoked by developer-written tests. Such a function is related to patches and not related to tests, and developers usually have specific expectations regarding its behavior. To identify target functions, we first instrument each patch-modified function in $R_{g}$ and $R_{o}$, and run all test files in $R_{g}$ and $R_{o}$ to collect call traces. Each call trace starts from a test function in $R_{g}$ or $R_{o}$ and ends at a patch-modified function. The first non-test function in each call trace satisfies the criterion mentioned above and PatchDiff annotates it as a target function. Different call traces can have the same target function. So a target function may correspond to multiple call traces.
@@ 3.2.2. Contextual Code Extraction
In each call trace, the functions in test files provide information about how to invoke the target function, and the functions in non-test files are either the target function or illustrate how the target function utilizes the patch-modified function. So all the functions in the call trace are useful for generating differentiating tests. For each target function, we extract the functions from the shortest call trace collected in $R_{g}$ and the shortest one collected in $R_{o}$. For each of these functions, we further map it to its before-patch version in $R_{t}$ based on its class and function names. The mapped functions provide the contextual information before patches are applied and are referred to as context functions. Only providing function definitions to the LLM may miss critical contextual information such as class information. Therefore, we extend context functions to construct contextual code. Specifically, we collect all Python files that contain the target function or at least one context/patch-modified function from $R_{t}$. We remove the functions that are not target, context, or patch-modified functions from these files. Any classes rendered empty after this deletion are also discarded. In addition, in test files, we annotate code lines where the target function is invoked with a comment to direct the LLM’s attention. All the streamlined files form the contextual code of this target function.
@@ 3.2.4. Unqualified Test Filtering
PatchDiff aims at generating tests for target functions to expose meaningful behavioral discrepancies. However, since LLMs do not always follow the instructions, the generated tests may not only examine the specified target function. To increase the probability of exposing meaningful behavioral discrepancies, PatchDiff filters out such tests. Specifically, we instrument patch-modified functions and execute each generated differentiating test to collect call traces. For each differentiating test, if there is a call trace where the function directly invoked by the test function is not a target function, we filter it out. Because this test does not only examine target functions. We further filter out flaky tests. In detail, for each remaining differentiating test, we run it under $P_{g}$ and $P_{o}$ for 20 times each. If it passes under one patch for all 20 times and fails under the other patch for at least one time, we regard it as valid. Otherwise, we filter out it. All the remaining differentiating tests are regarded as the output of PatchDiff.
@@ 4.2. RQ2: Revealing Behavioral Discrepancies Between Plausible and Oracle Patches
Table 3. API costs of PatchDiff under different repair iteration bounds
Max Repair Iter #Susp. Patches Cost/Patch(\$) Total Cost (\$)
2 260/877 (29.6%) 0.105 91.716
1 250/877 (28.5%) 0.075 66.186
0 228/877 (26.0%) 0.039 33.962
Table 3 details the API costs of PatchDiff. To get more insightful results in the empirical study, PatchDiff repairs the failed tests for up to 2 times, leading to a cost of 0.105\$ per patch. A lower repair iteration bound can reduce the cost to a minimum of 0.039\$ per patch, while only introducing an acceptable decrease (3.6%) of suspicious patches. These results indicate that PatchDiff is both effective and cost-efficient for large-scale use.
Table 4. Detected suspicious patches using different underlying LLMs (sampled 100 patches for each tool)
Model CodeStory LearnByInteract OpenHands Overall
GPT-4o-mini 26 (26.0%) 32 (32.0%) 26 (26.0%) 84 (28.0%)
DeepSeek-V3 34 (34.0%) 43 (43.0%) 40 (40.0%) 117 (39.0%)
Qwen3-A22B 45 (45.0%) 49 (49.0%) 49 (49.0%) 143 (47.7%)
4.2.2. Evaluating PatchDiff with Alternative LLMs.
To further demonstrate that PatchDiff is not tied to a specific underlying LLM, we also evaluate its performance with two open-source models, deepseek-v3 and qwen3-235b-a22b-instruct-2507. For this purpose, we randomly sample 100 plausible patches from each of the three evaluated tools and apply PatchDiff using these models.
Results:Table 4 presents the results. When powered by more capable LLMs, PatchDiff achieves substantially higher detection rates of suspicious patches, showing that its effectiveness generalizes across different underlying models. This confirms that PatchDiff is not coupled with GPT-4o-mini and can work with open-source models. Nevertheless, for the main study, we adopt GPT-4o-mini as the underlying model because it offers both low cost and high usability.
@@ 4.2.3. Comparison with Existing Test Generators
Approach: We use two representative test generators for Python (i.e., Pynguin (Lukasczyk and Fraser, 2022) and CoverUp (Pizzorno and Berger, 2024)), along with an issue reproduction tool LIBRO (Kang et al., 2024). Pynguin is the state-of-the-art search-based test generation tool for Python. CoverUp is a representative LLM-based test generation tool with its implementation publicly available and easy to use. Please note that both Pynguin and CoverUp support only Python 3.10 or higher, a requirement met by only 14.8% of the 500 tasks in SWE-bench Verified. Consequently, our evaluation is confined to the plausible patches generated by the three tools that are under a compatible Python version, yielding 133 plausible patches. For both test generators, we designate the patch-modified files as the target modules, and tests are generated on the repository with the oracle patch applied. Since CoverUp supports only PyTest and fails on the instances employing other test frameworks (e.g., Django’s customized framework), we do not provide original tests to CoverUp to ensure it can handle the 133 instances. Since LIBRO was originally implemented in Java, we re-implemented it in Python. We omit its selection phase, as it only ranks tests for developers and does not potentially increase differentiating tests; instead, we execute all generated tests to check for behavioral discrepancies between each plausible patch and its oracle. We use the best-performing configuration from the original paper (two examples, n=50n=50(Kang et al., 2024). For both CoverUp and LIBRO, we follow PatchDiff ’s setting and adopt gpt-4o-mini-2024-07-18 as the underlying LLM for fair comparison.
@@
Results: Table  5 presents the evaluation results for Pynguin and CoverUp. Pynguin produces at least one test file for only 3 plausible patches because it frequently encounters exceptions during generation. For example, for 93 patches, Pynguin raises an AttributeError which states that in the file pynguin/testcase/execution.py, attribute IN cannot be found in enum Compare, which may be a bug in Pynguin. CoverUp produces at least one test file only for 40 patches due to the difficulties it has in constructing appropriate execution environments (such as the database setup required by Django). None of the tests generated by Pynguin or CoverUp reveal behavioral discrepancies between the plausible and oracle patches, i.e., none of these tests are differentiating tests. In contrast, PatchDiff produces at least one test file for 117 patches and successfully generates differentiating tests for 56 of them. Note that among the 133 plausible patches, PatchDiff only attempts to generate tests for 117 of them. 15 of them are identical to their corresponding oracle patches and are omitted by PatchDiff, and 1 patch contains syntax errors and cannot be correctly parsed by the unidiff library (uni, 2025). Table 6 summarizes the evaluation results for LIBRO. Even under its best-performing configuration, LIBRO is able to generate differentiating tests for only 5.9% of the plausible patches, in contrast to 29.6% achieved by PatchDiff. This limited performance can be attributed to LIBRO’s reliance solely on issue descriptions for test generation, without considering the actual patch content. By contrast, PatchDiff explicitly targets the patch-relevant code and aims to generate tests that expose subtle behavioral differences between patches. These results underscore the effectiveness and necessity of PatchDiff in generating differentiating tests.
Table 6. Suspicious patches detected by PatchDiff and LIBRO
Tool CodeStory LearnByInteract OpenHands Overall
PatchDiff 91 (29.3%) 97 (32.2%) 72 (27.2%) 260 (29.6%)
LIBRO 18 (5.8%) 20 (6.6%) 14 (5.3%) 52 (5.9%)
Answers to RQ2: PatchDiff generates differentiating tests for 29.6% of the plausible patches. This suggests that a significant proportion of plausible patches behave differently from the oracle patches. In contrast, test generation baselines can only generate differentiating tests for up to 5.9% of the plausible pathces.